Tamarisk Corporation purchased a computer on December 31, 2019, for $126,000, paying $36,000 down and agreeing to pay the balance in five equal installments of $18,000 payable each December 31 beginning in 2020. An assumed interest rate of 8% is implicit in the purchase price. Prepare the journal entry at the date of purchase. Prepare the journal entry at December 31, 2020, to record the payment and interest (effective-interest method employed). Prepare the journal entry at December 31, 2021, to record the payment and interest (effective-interest method employed).
Answer #1

1.

Date Account Title and Explanation Debit Credit
December 31 2019 Equipment 93495
Discount on Notes Payable 32505
Notes Payable 90000
Cash 36000
(Recorded Purchase of Equipment)

2)

Date Account Title and Explanation Debit Credit
December 31 2020 Interest Expenses 4600
Notes Payable 18000
Discount on Notes Payable 4600
Cash 18000
(Recorded Payment of 1st installment & interest expenses)

3)

Date Account Title and Explanation Debit Credit
December 31 2021 Interest Expenses 3528
Notes Payable 18000
Discount on Notes Payable 3528
Cash 18000
(Recorded Payment of 2nd installment & interest expenses)

Working

1)Notes Payable = 18000*5 = 90000

Equipment = 36000 + 18000*(1-(1+8%)^-5)/10%

Equipment = 93495

Discount on Notes Payable = 126000-93495 = 32505

2) Interest Expense = (90000-32505)*8% = 4600

Amortization of Discount = 4600

Unamortized discount = 32505 - 4600= 27905

3) Interest Expenses = (90000-18000 - 27905)*8% = 3528